# Surajit Das The code will be provided upon request to the corresponding author.11 0009-0008-6692-6697    Pavel Zun 22 1111-2222-3333-$44$

### v1

@@ Conclusion
In summary, this software tool provides a unified platform for the entire image analysis workflow, from initial image inspection and interactive data exploration to advanced, configurable segmentation and quantitative morphological analysis. Its design is particularly suited for biomedical research applications requiring detailed characterization of cellular structures and spatial relationships.
{credits}
4.0.1 Acknowledgements
The research was supported by ITMO University Research Projects in AI Initiative (RPAII) (project #640103, Development of methods for automated processing and analysis of optical and atomic force microscopy images using machine learning techniques)*.
[…]
The authors declare that they have no competing interests relevant to the content of this article.acknowledgments [^4] This preprint is not peer-reviewed and is provided under a specific license. Unauthorized use is prohibited and is the sole responsibility of the user.,

### v5

@@ Conclusion
The proposed software system democratizes advanced microscopy image analysis by offering a low-cost, CPU-based, and training-free solution for label-free cytometric segmentation and auto-analysis. It enables researchers—particularly in low-resource biomedical laboratories—to perform accurate and reproducible cell segmentation, feature extraction, and statistical reporting without the need for GPU hardware or programming skills.
The framework integrates four core components: (1) semantic and instance segmentation using fuzzy logic and spatial statistics; (2) post-segmentation refinement via morphological operations; (3) quantitative morphometric analysis for geometric and spatial characterization of cells; and (4) automated report generation using an interactive profiler. The unified pipeline supports both GUI and command-line execution, enhancing accessibility across user levels.
Compared with tools such as CellProfiler, Cellpose, and StarDist, the proposed system provides a fully local, unsupervised alternative that removes dependence on training data, manual annotation, and high-performance computing. While deep learning approaches yield strong accuracy, their reliance on labeled datasets and GPUs restricts usability. In contrast, our CPU-based system achieves comparable precision and reproducibility on brightfield images of unstained cells.
By bridging the gap between algorithmic research and practical deployment, this software delivers a reproducible and interpretable platform for quantitative microscopic image analysis. Its adaptability to diverse morphologies and imaging modalities, along with automated reporting, put its value for biomedical research, including applications in cell transplantation, regenerative medicine, and tissue characterization, where precise structural & spatial quantification are essential [^5] This preprint is not peer-reviewed and is provided under a specific license. Unauthorized use is prohibited and is the sole responsibility of the user..
Disclosure of Interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Ethics Statement
This study did not involve human participants or animals.
Author Contributions
S.D. designed the study, analyzed the data, implemented the model, drafted the manuscript, and supervised the work. P.Z. provided supervision and critical review of the manuscript. All authors read and approved the final version